3.Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form
Answer by nyc_function(2741)

The equation we seek has the same slope as the given equation because we know they are parallel. We use the point-slope formula, the given point (-4,2) and given slope -1. Do you see in your equation -x? There is an invisible -1 next to x.
So, -x is the same as -1x.
The coefficient of x ( the number next to x) is the slope of the line.
y - (y coordinate of your point) = m(x -(x coordinate of your point)
Let m = slope of the line
y - 2 = -1(x -(-4))
y - 2 = -1(x + 4)
y - 2 = -x - 4
y = -x - 4 + 2
ANSWER: y = -x - 2